OpenStudy (anonymous):

Store Y sells two brands of socks, one at $3 a pair and the other at $4 a pair. If Janine must spend exactly $29 on socks and there is no sales tax, how many pairs of socks can she buy? a. 6 b. 7 c. 8 d. 9 e. 10

OpenStudy (khalilforthewin):

what do you think it is

OpenStudy (anonymous):

i don't know which is why i'm asking

OpenStudy (anonymous):

i dont need the answer i need to know how to do it

OpenStudy (khalilforthewin):

ok you will count times three until you cant go anymore then back track in fours as in 3,6,9,12,15,18,21,24,27 but you need two so you will go back so it stops at 21 then add 4 25 then add 4 again to get 29 and it took you 9 numbers to get to it so your answer would be?

OpenStudy (anonymous):

what do u mean need two

OpenStudy (anonymous):

Let’s say that Janine buys x of the first pair of socks and y of the second. This means that she spends 3x on the first and 4y on the second. You can rephrase the question as follows: if 3x + 4y = 29, what could be the value of x + y? (Keep in mind that x and y must be integers.)Consider all multiples of 4 that are less than 29, and find which of those multiples, when subtracted from 29, will leave you with a multiple of 3. Here are all the multiples of 4 under 29: 4, 8, 12, 16, 20, 24, and 28. Only 8 and 20 can leave you with a multiple of 3 when subtracted from 29. If Janine spends $8 on the second brand of socks, she buys 2 pairs. She then has $21 to spend on the first brand of socks, meaning that she can buy 7 pairs of the first brand. If you add up the pairs of socks, she buys 2 + 7 = 9 pairs. If she spends $20 on the second brand of socks, which gets her 5 pairs, then she has $9 left to spend on the first brand, at $3 per pair, which lets her buy 3 pairs. In this scenario she can buy 3 + 5 = 8 pairs. The correct answers are 8 and 9 pairs.